# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@ B.4 Proof of Proposition 2
Besides the conditional covariance matrix, our structural STVAR model defined in Equations (3.2) and (3.3) is obtained from the stochastic process $z_{t}$ defined in Equation (17) (or equally (15)) of Saikkonen (2008) as follows. First, we assume that the number of unit roots in the model is zero, so $n=r$ in Saikkonen (2008). Consequently, $z_{t}=y_{t}$ and, in the definition of the matrix $J$ defined above Equation (17) of Saikkonen (2008), $\beta=c=I_{d}$ (see Saikkonen, 2008, Note 2). The matrix $J$ is thereby nonsingular, and following Kheifets and Saikkonen (2020), we can choose the matrix $S$ in Equation (17) to be the inverse of $J^{\prime}$. Second, we assume the transition functions $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})$ of Saikkonen (2008) are independent of the random variable $\eta_{t}$. Hence, the transition functions are functions of $Z_{t-1}=(y_{t-1},...,y_{t-p})$, and we can write $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})=\alpha_{m,t}$. Using the notation of this paper, the first term on the right side of Equation (17) in Saikkonen (2008) then reduces to $\sum_{m=1}^{M}\alpha_{m,t}\sum_{i=1}^{p}A_{m,i}y_{t-i}$. Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}$ (with the notation of this paper).
@@
Secondly, the conditional covariance matrix appears in the proof of Saikkonen (2008) on page 313, where Saikkonen (2008) shows that an appropriate version of Condition (15.3) of Meyn and Tweedie (1993, p. 358) is satisfied. To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$, where $j$ is a positive integer, is bounded from above by a finite real number.
@@ C.1 Nonlinear least squares estimation
where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters. [^3] With threshold weights (2.3), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.4) weights, $\alpha=(c,\gamma)$. If the transition weights are exogenous, we simply treat them as known constants. The NLS estimates of the parameters $\phi_{m,0},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
@@
If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ they depend on needs to be estimated. Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only. However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that not able escape from local maxima are not sufficient. Since we only need approximate estimates to serve as starting values for Phase 2 and 3 estimation in the three-phase procedure described in Appendix 4.3, the following procedure is adopted.
@@
First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model. As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model and its parameters are not identified. Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
@@
One way to asses the contribution of each regime is by summing over the transition weights for each regime, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$. In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime. Therefore, one way to determine whether a regime contributes enough to the model is to check whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime for all $m=1,...,M$. Denote the large enough sum as $\tilde{T}$. We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough. [^4] Our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
@@
Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range. For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number some large positive number that implies almost discrete regime-switches. With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$. Hence, it is advisable to select a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time. Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$. Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$. Denote this subset as $\mathcal{W}$.
@@
Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1,0},....,\phi_{M,0}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$. Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
@@ C.2 Penalized nonlinear least squares estimation
where the penalization term $P(\alpha)$ increases as the parameter values approach the boundary of, or move further into, an uninteresting region of the parameter space. Here we are interested only in avoiding estimates that do not satisfy the usual stability condition for all the regimes. Therefore, we define the penalty term as
@@
The penalized NLS estimates of $\phi_{m,0},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2). In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
@@ Appendix D A Monte Carlo study
This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-phase estimation procedure discussed in Section 4. We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable. To maintain some constancy, we use the same impact matrices across all specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$. For TVAR models, we use the threshold parameter value $r_{1}=0.8$ and for LSTVAR models the location and scale parameter values $c=0.8$ and $\gamma=5.0$.
For both the TVAR model and the LSTVAR model, we consider two specifications of the AR matrices, presented in Table D.1. In the first specification (labelled as TVAR 1 and LSTVAR 1 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the stability region (see Section 4.1). In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region. To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus). The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specification (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
@@
Table D.1: The parameters values of $\phi_{1,0},\phi_{2,0},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study. Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m,0}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$. Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-phase estimation procedure based on the penalized log-likelihood function, proposed in Section 4. After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values. To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
@@
Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-phase estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-phase estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
First, the estimator of the second degrees-of-freedom parameter (true parameter value $\nu_{2}=12$) is substantially biased upwards in samples of length $T=250,500$, and $1000$, and also the standard deviation of the estimates is very large. This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution. Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values, and hence, possibly very large degrees-of-freedom parameter values drawn by random in Phase 2 estimation may remain large through Phase 3 estimation. Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the upward bias should not be a concern (for example, if only the Monte Carlo repetitions in which the estimate of $\nu_{2}$ is less than $100$ are considered, then for the specification TVAR 1, the estimated bias reduces to $5.37$ and the standard deviation to $17.63$ when $T=250$).
@@
Third, somewhat surprisingly, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small. A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters. In particular, the estimates may be sensitive to extreme values drawn for the skewed $t$ shocks, potentially amplifying small-sample bias.

### v4

@@ B.4 Proof of Proposition 2
Besides the conditional covariance matrix, our structural STVAR model defined in Equations (3.2) and (3.3) is obtained from the stochastic process $z_{t}$ defined in Equation (17) (or equally (15)) of Saikkonen (2008) as follows ($z_{t}$ of Saikkonen (2008) should not be confused with the switching variable of this paper). First, we assume that the number of unit roots in the model is zero, so $n=r$ in Saikkonen (2008). Consequently, $z_{t}=y_{t}$ and, in the definition of the matrix $J$ defined above Equation (17) of Saikkonen (2008), $\beta=c=I_{d}$ (see Saikkonen, 2008, Note 2). The matrix $J$ is thereby nonsingular, and following Kheifets and Saikkonen (2020), we can choose the matrix $S$ in Equation (17) to be the inverse of $J^{\prime}$. Second, we assume the transition functions $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})$ of Saikkonen (2008) are independent of the random variable $\eta_{t}$. Hence, the transition functions are functions of $Z_{t-1}=(y_{t-1},...,y_{t-p})$, and we can write $h_{s}(S^{\prime}JZ_{t-1},\eta_{t})=\alpha_{m,t}$. Using the notation of this paper, the first term on the right side of Equation (17) in Saikkonen (2008) then reduces to $\sum_{m=1}^{M}\alpha_{m,t}\sum_{i=1}^{p}A_{m,i}y_{t-i}$. Third, assume the set $\text{l}_{2}$ is empty, which implies that the second term on the right side of Equation (17), defined in Equation (11) of Saikkonen (2008), reduces to $\sum_{m=1}^{M}\alpha_{m,t}\phi_{m}$ (with the notation of this paper).
@@
Secondly, the conditional covariance matrix appears in the proof of Saikkonen (2008) on page 313, where Saikkonen (2008) shows that an appropriate version of Condition (15.3) of Meyn and Tweedie (1993, p. 358) is satisfied. To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$ is bounded from above by a finite real number, where $j$ is a positive integer, $\mathcal{A}=\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\}$, and $\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}$ as well as the set $\mathcal{A}^{j}$ are defined in Appendix A.
@@ C.1 Nonlinear least squares estimation
where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters. [^5] With threshold weights (2.3), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.4) weights, $\alpha=(c,\gamma)$. If the transition weights are exogenous, we simply treat them as known constants. The NLS estimates of the parameters $\phi_{m},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
@@
If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ (they depend on) needs to be estimated. Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only. However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are not sufficient. Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Appendix 4.3, the following procedure is adopted.
@@
First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model. As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model at all. Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
@@
One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$. In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime. Therefore, one way to determine whether a regime contributes enough to the model is to check whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime for all $m=1,...,M$. Denote the large enough sum as $\tilde{T}$. We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough. [^6] We acknowledge that our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
@@
Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range. For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number some large positive number that implies almost discrete regime-switches. With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$. Hence, we recommend selecting a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time. Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$. Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$. Denote this subset as $\mathcal{W}$.
@@
Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1},....,\phi_{M}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$. Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
@@ C.2 Penalized nonlinear least squares estimation
where the penalization term $P(\alpha)$ increases as the parameter values approach the boundary of, or move further into, an uninteresting region of the parameter space. To focus on avoiding estimates that do not satisfy the usual stability condition for all regimes, we define the penalty term as
@@
The penalized NLS estimates of $\phi_{m},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2). In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
@@ Appendix D A Monte Carlo study
This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-step estimation procedure discussed in Section 4. We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable. To maintain some constancy, we use the same impact matrices across all specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$. For TVAR models, we use the threshold parameter value $r_{1}=0.8$ and for LSTVAR models the location and scale parameter values $c=0.8$ and $\gamma=5.0$.
For both the TVAR model and the LSTVAR model, we consider two specifications of the AR matrices, presented in Table D.1. In the first specification (labelled as TVAR 1 and LSTVAR 1 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the stability region (see Section 4.1). In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region. To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus). The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
@@
Table D.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study. Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$. Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4. After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values. To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
@@
Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
First, the estimator of the second degrees-of-freedom parameter (true parameter value $\nu_{2}=12$) is substantially biased upwards in samples of length $T=250,500$, and $1000$, and also the standard deviation of the estimates is very large. This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution. Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values. Hence, possibly very large degrees-of-freedom parameter values drawn by random in Step 2 estimation may remain large through Step 3 estimation. Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the large upward bias should not be a concern.
@@
Third, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small. A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters. In particular, the estimates may be sensitive to extreme values drawn for the heavy-tailed skewed $t$ shocks, potentially amplifying small-sample bias.